# MCM: Multi-layer Concept Map for Efficient Concept Learning from Masked Images

### v1

@@ 1 Introduction
(1) We propose the Multi-layer Concept Map (MCM) method to facilitate masked concept learning through vision reconstruction tasks, which involves the masked concept encoder and multi-layer concept mapping architecture (Section 3.1).
(2) This study investigates two dedicated loss functions to enhance the model’s ability in concept prediction, especially for training on unbalanced concept classes, i.e., the disentanglement loss and weighted concept loss (Section 3.1.2).
(3) The extensive quantitative and qualitative analysis involves concept prediction performance, reconstruction quality measured by Fréchet Inception Distance, computational cost, and diverse visualizations. MCM learns effective concept tokens using less than 75% of image patches while achieving competitive performance (Section 4.2, 4.3).
@@ 2.1 Masked image reconstruction
Masked image modeling has emerged as a pivotal learning technique in computer vision He et al. (2022); Yue et al. (2023); Zheng et al. (2023); Chen et al. (2023); Fu et al. (2024). For instance, Masked Autoencoders He et al. (2022) learn to reconstruct missing patches given only a small subset of visible patches reducing computational cost. Masked Diffusion Transformer Zheng et al. (2023) demonstrates enhanced training efficiency and generation results using a denoising diffusion objective on masked images. Cross-Attention Masked Autoencoders Fu et al. (2024) used cross-attention in the decoder to query visible tokens for masked image patch reconstruction. The cross-attention component takes a weighted sum of the visible tokens across different input blocks to fuse the features for each decoder block, leveraging low-level information for reconstruction.
@@ 2.2 Disentangled representation learning
Disentangled concept learning Bengio et al. (2013); Higgins et al. (2017); Locatello et al. (2019); Härkönen et al. (2020); Anirudh et al. (2021); Yang et al. (2022); Sun et al. (2023); Ismail et al. (2023) aims to uncover the underlying explanatory factors hidden within observed data. For example, methods such as -VAE Higgins et al. (2017) and FactorVAE Kim & Mnih (2018) search for directions in the latent space that correlate with distinct human-interpretable concepts. Moreover, Concept Tokenization Yang et al. (2022) focuses on learning disentangled object representations and inspecting latent traversals for various factors. Additionally, Concept Bottleneck models Ismail et al. (2023); Yuksekgonul et al. (2022); Oikarinen et al. (2023); Yang et al. (2023) learn representations that correspond to specific human-understandable concepts. Energy-based methods Du et al. (2020); Li et al. (2022) aim to compute energy functions of various concepts and combine their probability distributions achieving conjunctions, disjunctions, and negations of various concepts.
@@ 3 Method
3.1 Multi-layer concept map for masked concept learning
3.1.1 Concept encoding from masked images
@@
3.1.2 Image decoding with multi-layer concept mapping
@@
Masked reconstruction loss.
@@
Weighted concept loss.
@@ 4.1.1 Datasets
We employ the CelebA Liu et al. (2015) dataset, a face attributes dataset characterized factors of variation, where we selected 11 concepts with varying frequencies (please refer to Appendix A.2), including ‘Bald,’ ‘Bangs,’ ‘Black Hair,’ ‘Blond Hair,’ ‘Chubby,’ ‘Eyeglasses,’ ‘Mustache,’ ‘Wearing Hat,’ ‘Male,’ ‘Smiling,’ and ‘Young.’ Antonyms were generated by appending ‘Not’ before each concept (e.g., ‘Smiling’ vs. ‘Not Smiling’).
@@ 4.2 Quantitative results
4.2.1 Concept prediction and reconstruction performance
We aim to study how various training mask ratios affect the model’s performance in the concept prediction and reconstruction tasks. We measure numerical prediction performance based on accuracy, precision, recall, and F1 score, evaluating reconstruction performance based on the Fréchet Inception Distance (FID) score. Table 1 demonstrates the impact of varying mask ratios, showing an optimized mask ratio of 0.5 for the small-sized model and 0.25 for the base and large-sized models. For each entry, we trained the small and base-sized models for 500 epochs and the large-sized model for 100 epochs. In particular, for the base and large-sized models, as the mask ratio increases beyond 0.25, accuracy and F1-score start to decline, and FID increases substantially.
| Model Size | Mask Ratio | Accuracy | Precision | Recall | F1-Score | FID | Training T (h) |
@@ 4.2.1 Concept prediction and reconstruction performance
4.2.2 Ablation studies
(2) W/O Learnable Latent Concepts: Concept label embeddings query encoder layers, with the weighted output serving as input to specific decoder layers.
(3) Repetitive Latent Concepts: The learned concept tokens in the latent space are distributed across all decoder layers.
@@ 4.2.2 Ablation studies
Comparison with Masked Autoencoder.
Table 3 demonstrates the comparison with the Masked Autoencoder (MAE) He et al. (2022) on the concept learning tasks. We follow the architecture designs and hyperparameter settings of a base-sized MAE. For MAE, we added a linear layer in its latent space to predict concepts using the visible tokens, training the model from scratch. Additionally, the binary concept labels from the CelebA dataset were employed for each concept position.
| Method | Accuracy | Precision | Recall | F1-Score | FID |
| MCM | 0.945 | 0.881 | 0.839 | 0.860 | 1.884 |
| Masked Autoencoder | 0.537 | 0.809 | 0.75 | 0.778 | 1.031 |
@@ 4.3 Qualitative results
Masked image reconstruction and editing.
Masked image reconstruction involves predicting the original image from a partially masked input, where patches are randomly removed. We demonstrate the performance of our approach (training mask ratio 25%) in reconstructing images with a significant portion of patches masked during testing (e.g., 75%) in Figure 2(a). Additionally, by providing specific text-based concept tokens, we can traverse the concept latent space and manipulate the reconstruction of the masked patches. The resulting image is reconstructed to align with both the contextual unmasked patches and the specified concepts. Notably, we can either activate each concept individually or combine multiple concepts, as illustrated in Figure 2(b). We also compare editing results across various model sizes with different computational costs, highlighting that the model’s ability to perform masked image editing and reconstruction improves progressively as it scales.
@@ A.1 Model settings
Common parameters across all models included the AdamW optimizer with a learning rate of , a weight decay of 0.01, four self-attention heads, hidden layer size of 512, and a batch size of 1024. The Small model featured 2 encoder layers, 1 decoder layer, an MLP size of 128, and was trained for 500 epochs on 48x48 images with a patch size of 6. The Base model expanded to 6 encoder layers, 3 decoder layers, an MLP size of 512, and used 96x96 images with a patch size of 8. Finally, the Large model utilized 12 encoder layers, 6 decoder layers, an MLP size of 1024, and 192x192 images with the same patch size of 8 but was trained for only 100 epochs due to its increased complexity. The coefficients and have a default value of one.
@@ A.2 Unbalanced concept class distribution in CelebA
A.3 Additional testing time results
| Mask Ratio | Accuracy | Precision | Recall | F1-Score | FID |
| 0.0 | 0.940 | 0.854 | 0.812 | 0.832 | 1.520 |
| 0.1 | 0.944 | 0.824 | 0.794 | 0.809 | 1.856 |
| 0.25 | 0.947 | 0.877 | 0.824 | 0.850 | 1.843 |
| 0.5 | 0.946 | 0.881 | 0.819 | 0.849 | 2.112 |
| 0.75 | 0.946 | 0.890 | 0.812 | 0.849 | 3.790 |
| 0.9 | 0.938 | 0.840 | 0.776 | 0.807 | 8.045 |
A.4 Design of the testing time mask

### v2

@@ 1 Introduction
(1) We propose the Multi-layer Concept Map (MCM) method to facilitate masked concept learning through vision reconstruction tasks, which involves the masked concept encoder and multi-layer concept mapping architecture.
(2) This study investigates two dedicated loss functions to enhance the model’s ability in concept prediction, especially for training on unbalanced concept classes, i.e., the disentanglement loss and weighted concept loss.
(3) The extensive quantitative and qualitative analysis involves concept prediction performance, reconstruction quality measured by Fréchet Inception Distance, computational cost, and diverse visualizations. MCM learns effective concept tokens using less than 75% of image patches while achieving competitive performance.
@@ 2.1 Masked image reconstruction
Masked image modeling has emerged as a pivotal learning technique in computer vision He et al. (2022); Yue et al. (2023); Zheng et al. (2023); Chen et al. (2023); Fu et al. (2024). For instance, Masked Autoencoders He et al. (2022) learn to reconstruct missing patches given only a small subset of visible patches reducing computational cost. Masked Diffusion Transformer Zheng et al. (2023); Gao et al. (2023) demonstrates enhanced training efficiency and generation results using a denoising diffusion objective on masked images. Cross-Attention Masked Autoencoders Fu et al. (2024) used cross-attention in the decoder to query visible tokens for masked image patch reconstruction. The cross-attention component takes a weighted sum of the visible tokens across different input blocks to fuse the features for each decoder block, leveraging low-level information for reconstruction.
@@ 2.2 Disentangled representation learning
Disentangled concept learning Bengio et al. (2013); Higgins et al. (2017); Locatello et al. (2019); Härkönen et al. (2020); Anirudh et al. (2021); Yang et al. (2022); Sun et al. (2023); Ismail et al. (2023) aims to uncover the underlying explanatory factors hidden within observed data. For example, methods such as -VAE Higgins et al. (2017) and FactorVAE Kim & Mnih (2018) search for directions in the latent space that correlate with distinct human-interpretable concepts. Moreover, Concept Tokenization Yang et al. (2022) focuses on learning disentangled object representations and inspecting latent traversals for various factors. Additionally, Concept Bottleneck models Ismail et al. (2023); Yuksekgonul et al. (2022); Oikarinen et al. (2023); Yang et al. (2023) learn representations that correspond to specific human-understandable concepts. Energy-based methods Du et al. (2020); Li et al. (2022); Du et al. (2023) aim to compute energy functions of various concepts and combine their probability distributions achieving conjunctions, disjunctions, and negations of various concepts.
@@ 3 Method
3.1 Concept encoding from masked images
@@
3.2 Image decoding with multi-layer concept mapping
@@
3.2.1 Masked reconstruction loss
@@ 3.2.2 Disentanglement loss
3.2.3 Weighted concept loss
@@ 4.1.1 Datasets
We conduct our experiments on the CelebA dataset Liu et al. (2015), a large-scale collection of celebrity face images annotated with rich binary attributes that capture diverse factors of variation. From this dataset, we select a subset of 11 semantically meaningful concepts exhibiting different levels of frequency and visual distinctiveness (please refer to Appendix A.2 for distribution details). To enable systematic evaluation across positive and negative attributes, we construct antonymic counterparts by prefixing each concept with “Not” (e.g., Smiling vs. Not Smiling). This design enables performance evaluation under both frequent and underrepresented concepts, while ensuring a balanced treatment of concept presence and absence.
@@ 4.2 Quantitative results
| Size | Mask Ratio | Accuracy | Precision | Recall | F1-Score | FID | Training T (h) |
@@
| Method | Accuracy | Precision | Recall | F1-Score | FID | Size (M) |
| MCM | 0.945 | 0.881 | 0.839 | 0.860 | 1.884 | 112.2 |
| MCM + Weighted and Disentangle Losses | 0.946 | 0.886 | 0.859 | 0.872 | 1.605 | |
| Masked Autoencoder He et al. (2022) | 0.537 | 0.809 | 0.750 | 0.778 | 1.274 | 114.8 |
| Masked DiT v2 Gao et al. (2023) | 0.735 | 0.189 | 0.241 | 0.146 | 32.38 | 87.0 |
| Cross-Attention MAE Fu et al. (2024) | 0.703 | 0.437 | 0.673 | 0.469 | 3.295 | 53.2 |
| Concept Tokenization Yang et al. (2022) | 0.925 | 0.756 | 0.686 | 0.719 | 3.971 | 118.7 |
| CLIP-Based MAE He et al. (2022) | 0.908 | 0.788 | 0.731 | 0.759 | 6.059 | 94.5 |
@@
4.2.1 Concept prediction and reconstruction performance
We aim to study how various training mask ratios affect the model’s performance in the concept prediction and reconstruction tasks. We measure numerical prediction performance based on accuracy, precision, recall, and F1 score, evaluating reconstruction performance based on the Fréchet Inception Distance (FID) score. Table 1 demonstrates the impact of varying training mask ratios, showing an optimized mask ratio of 0.5 for the small-sized model and 0.25 for the base and large-sized models. For each entry, we trained the small and base-sized models for 500 epochs and the large-sized model for 100 epochs. In particular, for the base and large-sized models, as the mask ratio increases beyond 0.25, accuracy and F1-score start to decline, and FID increases substantially.
Table 2 demonstrates the comparison with baseline methods including the Masked Autoencoder (MAE) He et al. (2022), the Masked Diffusion Transformer v2 (Masked DiT v2) Gao et al. (2023), the Cross-Attention MAE Fu et al. (2024), the Concept Tokenization method Yang et al. (2022), and the CLIP-Based MAE He et al. (2022). For the CLIP-Based MAE model, we employ concept embeddings from the pretrained CLIP model to supervise representation learning in the MAE’s latent space. For the other methods, we follow the architecture designs and hyperparameter settings of the base-sized model. For concept prediction, we leverage the latent representations of the unmasked visible tokens as input to an MLP attached to the model, with binary concept labels from the CelebA dataset at each concept position. All models are trained from scratch. Note that our method (MCM) does not require binary concept labels for training. Instead, it learns a set of specific concept tokens directly in the latent space, as illustrated in Figure 2.
The comparison demonstrates that, even with supervised guidance from binary labels, MAE falls short of achieving concept-learning performance comparable to the proposed MCM method. The advantages of the proposed MCM are evident not only in its superior concept prediction accuracy but also in its novel image editing capabilities for masked image reconstruction, which is a functionality that MAE cannot provide. Additional comparison of the reconstruction results are presented in Figure 6. Moreover, inducing the concept-learning objective into other methods, such as Masked DiT v2, substantially degrades their masked reconstruction ability, as reflected in lower FID scores. In comparison to MCM, these methods neither attain competitive concept-learning performance nor support editing of masked images.
4.2.2 Ablation studies
(2) W/O Learnable Latent Concepts: We replace the learnable concept tokens at each encoder layer with the fixed concept label embeddings , which query the encoder layer’s output via cross-attention. The resulting weighted encoder output is then employed as the input to the corresponding decoder layer, i.e., .
(3) Repetitive Latent Concepts: The learned concept tokens in the latent space are used as the keys and values for cross-attention computations across all decoder layers. In other words, we replace the encoder’s layer-wise learned concept embeddings with the same latent concept embeddings.
@@ 4.3 Qualitative results
4.3.1 Masked image reconstruction and editing
Masked image reconstruction involves predicting the original image from a partially masked input, where patches are randomly removed. We demonstrate the performance of our approach (training mask ratio 25%) in reconstructing images with a significant portion of patches masked during testing (e.g., 75%) in Figure 3(a). Additionally, by providing specific text-based concept tokens, we can traverse the concept latent space and manipulate the reconstruction of the masked patches. The resulting image is reconstructed to align with both the contextual unmasked patches and the specified concepts. Notably, we can either activate each concept individually or combine multiple concepts, as illustrated in Figure 3(b). We also compare editing results across various model sizes with different computational costs, highlighting that the model’s ability to perform masked image editing and reconstruction improves progressively as it scales. In comparison to MCM, methods such as Masked Autoencoder He et al. (2022) lack the ability for image editing while reconstructing the masked image patches. Inducing the concept learning ability in methods such as Masked DiT v2 Gao et al. (2023) can also lead to a substantial degradation in output image quality, as shown in Appendix A.4.
@@ A.1 Model settings
Common parameters across all models (Table 4) include the AdamW optimizer with a learning rate of , a weight decay of 0.01, four self-attention heads, hidden layer size of 512, and a batch size of 1024. The Small model features 2 encoder layers, 1 decoder layer, an MLP size of 128, and was trained for 500 epochs on 48x48 images with a patch size of 6. The Base model expands to 6 encoder layers, 3 decoder layers, an MLP size of 512, and used 96x96 images with a patch size of 8. Finally, the Large model utilizes 12 encoder layers, 6 decoder layers, an MLP size of 1024, and 192x192 images with the same patch size of 8 but was trained for only 100 epochs due to its increased complexity. The coefficients and have a default value of one.
@@ A.2 Unbalanced concept class distribution in CelebA
A.3 Algorithm
The complete algorithm is demonstrated in Algorithm 1.
A.4 Comparison of masked image reconstruction quality
We demonstrate in Figure 6 the difference in reconstruction quality of the methods discussed in the experiment section. In particular, we found that although the Masked Autoencoder (MAE) achieves a better FID score, its reconstruction results cannot be edited using the learned concepts. Moreover, methods such as MDTv2 suffer from substantial degradation in output image quality. Our method MCM, however, achieved a balance between reconstruction quality and concept-based image editing ability.
A.5 Design of the testing time mask